Schedule of cash payments for a service company

Horizon Financial Inc. was organized on February 28. Projected selling and administrative expenses for each of the first three months of operations are as follows:

March $82,200
April 77,300
May 70,300

Depreciation, insurance, and property taxes represent $18,000 of the estimated monthly expenses. The annual insurance premium was paid on February 28, and property taxes for the year will be paid in June. 72% of the remainder of the expenses are expected to be paid in the month in which they are incurred, with the balance to be paid in the following month.

Prepare a schedule of cash payments for selling and administrative expenses for March, April, and May.

Excel Learning Systems Inc.
Schedule of Cash Payments for Selling and Administrative Expenses
For the Three Months Ending May 31
March April May
March expenses:
Paid in March $
Paid in April $
April expenses:
Paid in April
Paid in May $
May expenses:
Paid in May
Total cash payments $ $ $
Answer #1

Answer-

HORIZON FINANCIAL INC.
SCHEDULE OF CASH PAYMENTS FOR SELLING AND ADMINISTRATIVE EXPENSES
FOR THE THREE MONTHS ENDING MAY 31
PARTICULARS MARCH APRIL MAY
$ $ $
March expenses:-
Paid in March ($82200-$18000)*72%=$46224
Paid in April ($82200-$18000)*28%=$17976
April expenses:-
Paid in April ($77300-$18000)*72%=$42696
Paid in May ($77300-$18000)*28%=$16604
May expenses:-
Paid in May ($70300-$18000)*72%=$37656
Total cash payments $ 46224 60672 54260